You are the operations manager for an airline and you are considering a higher fare level for passengers in aisle seats. How many randomly selected air passengers must you​ survey ? Assume that a prior survey suggests that about 73​% of air passengers prefer an aisle seat. Assume that you want to be 97% confident that the sample percentage is within 2.5 percentage points of the true population percentage.
Answer #1

Solution:

Given,

E = 0.025

c = 97% = 0.97

\hat p or p = 73% = 0.73

1 - \hat p 1- p = 1 - 0.73 = 0.27

Now,

\alpha = 1 - c = 1 - 0.97 = 0.03

\alpha/2 = 0.015

\thereforea/2 = 2.17 (using z table)

The sample size for estimating the proportion is given by

n = /2P(1 - p) E2

= (2.17)2 * 0.73 * 0.27 / (0.0252)

= 1484.9987

= 1485 ..(round to the next whole number)

Answer : Required Sample size is n = 1485

1485 randomly selected air passengers must you​ survey.